OpenStudy (anonymous):

Mary drove from A to B at a speed of 50 mi/h. On the way back, she drove at 60 mi/h. The total trip took 4 2/5 h of driving time. Find the distance between these two cities.

OpenStudy (anonymous):

Let d be the distance between A to B. r*t = d, t = d/r Solve the following for d:\[\frac{d}{50}+\frac{d}{60}=4+\frac{2}{5} \]d = 120 miles